if the degree measure of the smallest angle of a triangle is one-half of the degree measure of the second largest angle and one-third of the degree measure of the largest angle, then what is the degree measure of each angle?
Please help me with this

Answer by venugopalramana(3286)

if the degree measure of the smallest angle of a triangle is one-half of the degree measure of the second largest angle and one-third of the degree measure of the largest angle, then what is the degree measure of each angle?
LET THE SMALLEST ANGLE =X DEGREES
THIS ONE HALF OF SECOND LARGEST ANGLE
HENCE SECOND LARGEST ANGLE =2X
SMALLEST ANGLE IS ONE THIRD OF THE LARGEST ANGLE.
HENCE LARGEST ANGLE =3X
TOTAL OF 3 ANGLES =X+2X+3X=6X..BUT TOTAL OF ALL 3 ANGLES IN A TRIANGLE =180 DEGREES
HENCE 6X = 180
X=180/6 = 30 DEGREES .HENCE THE 3 ANGLES ARE 30,60 AND 90 DEGREES.